### v1

# SAT-based Exact Modulo Scheduling Mapping for Resource-Constrained CGRAs

Abstract.
Coarse-Grain Reconfigurable Arrays (CGRAs) represent emerging low-power architectures designed to accelerate Compute-Intensive Loops (CILs). The effectiveness of CGRAs in providing acceleration relies on the quality of mapping: how efficiently the CIL is compiled onto the platform. State of the Art (SoA) compilation techniques utilize modulo scheduling to minimize the Iteration Interval (II) and use graph algorithms like Max-Clique Enumeration to address mapping challenges. Our work approaches the mapping problem through a satisfiability (SAT) formulation. We introduce the Kernel Mobility Schedule (KMS), an ad-hoc schedule used with the Data Flow Graph and CGRA architectural information to generate Boolean statements that, when satisfied, yield a valid mapping. Experimental results demonstrate SAT-MapIt outperforming SoA alternatives in almost 50% of explored benchmarks. Additionally, we evaluated the mapping results in a synthesizable CGRA design and emphasized the run-time metrics trends, i.e. energy efficiency and latency, across different CILs and CGRA sizes. We show that a hardware-agnostic analysis performed on compiler-level metrics can optimally prune the architectural design space, while still retaining Pareto-optimal configurations. Moreover, by exploring how implementation details impact cost and performance on real hardware, we highlight the importance of holistic software-to-hardware mapping flows, as the one presented herein.
Coarse-Grained Reconfigurable Arrays, Hardware acceleration, Modulo Scheduling
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Hardware Hardware acceleratorsccs: Computer systems organization Reconfigurable computingccs: Software and its engineering Compilers
1. Introduction
The constant growth of computational requirements in everyday applications has increased the demand for high-performance and low-power architectures. Such architectures need to perform compute-intensive tasks efficiently, while at the same time dealing with tight power/resource constraints.
While Application Specific Integrated Circuits (ASICs) accelerators have been largely adopted in these scenarios due to their efficiency, they are limited by their fixed functionality and their lengthy and costly design time. The reconfigurability of Field Programmable Gate Arrays (FPGAs) can address these limitations, but at the cost of lower power efficiency compared to ASICs, due to their fine-grained structure.
Figure 1. Overview of the CGRA scheduling and mapping workflow. a): A CIL is identified in the source C code, b): Its associated DFG is derived by compiler analysis. c): An abstract view of available hardware resources (i.e., number of processing elements and their connectivity) is constructed. d): The nodes of the DFG are mapped on the processing elements while complying with architectural constraints. e): Nodes are modulo scheduled across different CGRA-instructions, partially overlapping the execution of multiple iterations. In the example, a new iteration is started at each CGRA-instruction, hence the scheduling has an Initiation Interval (II) equal to 1.
Coarse-Grain Reconfigurable Arrays (CGRAs) provide a meet-in-the-middle approach, providing run-time reconfigurability at the level of arithmetic operations, with very high computational efficiency (Li et al., 2021; Karunaratne et al., 2017). These characteristics are well suited in domains such as streaming and multimedia applications (Li et al., 2021; Akbari et al., 2018; Karunaratne et al., 2017; Oh et al., 2009; Lee et al., 2015; Wijerathne et al., 2019; Duch et al., 2017), but also in the edge domain, where they have been shown to enable flexible hardware acceleration in resource-constrained scenarios.
A CGRA is a mesh of Processing Elements (PEs) organized in a two-dimensional grid; each PE contains an Arithmetic-Logic Unit (ALU) and a number of internal registers. PEs are connected to their neighbors according to a mesh topology. In addition, shared connections are usually present from PEs to external memory to load inputs and store outputs. Computation in CGRAs is organized in CGRA-cycles, during which every PE performs an operation. The list of supported operations defines the Instruction Set Architecture (ISA) of a CGRA.
One of the fundamental challenges of CGRA exploitation is the compilation process, i.e. the translation of high-level source code onto CGRA code, while taking advantage of the parallelism offered by the architecture. To do so, a technique called modulo scheduling is traditionally employed, which constructs a Data Flow Graph (DFG) from a Compute-Intensive Loop (CIL), and then maps DFG nodes onto architecture PEs in an interleaved manner so that nodes from different iterations coexists in the same CGRA-cycle, minimizing the overall latency of each iteration. Figure 1 portrays an overview of this process, where a section of C code is extracted and mapped into a $2\times 2$ CGRA. First, a loop is marked for CGRA acceleration. Then, its DFG, which illustrates operations as nodes and data dependencies as edges, is derived. The DFG is then scheduled over the PEs of the target architecture, at different CGRA-cycles, while abiding to resource and dependency constraints. A possible modulo scheduling of the simple DFG in Figure 1.b is shown in Figure 1.e, highlighting the interleaving among loop iterations.
As the number of PEs (and the resources available to them) is reduced to fit on edge devices, mapping tools are faced with growing constraints. Existing techniques for modulo scheduling, described in Section 2, rely mainly on heuristics to schedule, place, and then route operations and data on the PEs. Here, we propose to address the mapping problem using SAT-MapIt, a satisfiability (SAT)-based formulation where data dependency, architectural constraints, and schedule are expressed as Boolean constraints. A valid mapping is then identified by determining if an assignment of the variables exists to satisfy the constructed Boolean formula. We show that this technique is able to efficiently explore the space of possible mappings better than State of the Art (SoA) techniques, hence producing high-performance mappings even for a tightly constrained CGRA topology.
We showcase the performance of our approach as part of a complete code-to-hardware framework. The flow starts from annotated C code and has at its endpoint a configuration bitstream that governs the execution of a (simulated or synthesized) CGRA. To this end, we herein target the open-hardware CGRA design in (Rodríguez Álvarez et al., 2023). We highlight that compiler-level metrics (such as Iteration Interval (II) and Utilization (U)) do indeed correlate with the run-time performance of a mapped CIL. However, hardware analysis (and hence the availability of a link between the compiler and the hardware) is the key to accurately assess the energy and timing requirements for a given CIL and CGRA architecture.
This paper is organized as follows. Section 2 reviews existing literature; Section 3 introduces foundational elements of our work and our problem formulation; Section 4 describes our SAT-based methodology and Section 5 compares the results obtained by SAT-MapIt with respect to SoA alternatives. Section 6 illustrates the developed code-to-hardware framework used to validate our approach targeting resource-constrained hardware while Section 7 shows execution measurements and analyzes the mapping and run-time performance metrics to prune the design space. Finally, Section 8 concludes the article.
2. Related Work
A recent survey (Podobas et al., 2020) has summarized the evolution of CGRA architectures and methodology advances in the last thirty years of research. Herein, we focus in particular on the CGRAs mapping problem, i.e. the compilation of software source code onto CGRA code, defined as the mapping of CILs instructions onto PEs. Existing methodologies can be divided into two main categories: the first using heuristics and the second using exact solutions.
Early works in the first category includes the method proposed by Mei et al.(Mei et al., 2007), which formulates scheduling, placement, and routing problems altogether, and proposes to solve them using simulated annealing. The attempt to solve all problems jointly does lead to long execution times, and potentially also to low quality mapping due to the difficulty to find good solutions in such a large space. Alternatively, in (Park et al., 2008) an edge-centric approach to modulo-scheduling was presented, where a schedule is generated by first routing each edge in the dataflow graph, and placement is addressed as a by-product of routing. In (Hamzeh et al., 2012), EPImap proposed to use both routing and re-computation to find a valid mapping, by adopting an epimorphic (time-extended) graph formulation enabling efficient identification of valid solutions. By adding nodes to the data dependency graphs, EPImap can find new valid solutions to the mapping while re-scheduling is performed to improve the efficiency. EPImap’s performance was later improved by GraphMinor (Chen and Mitra, 2014) and REGIMap (Hamzeh et al., 2013), by reducing the mapping problem to the graph minor and max clique problem. In turn, (Dave et al., 2018) RAMP authors have further refined REGIMap by explicitly modeling and exploring various routing strategies and choosing the best one for each given CIL. CRIMSON(Balasubramanian and Shrivastava, 2020) then proposed a randomized Iterative Modulo Scheduling (IMS) algorithm that explored the scheduling space more efficiently, and PathSeeker (Balasubramanian and Shrivastava, 2022) improved on CRIMSON (Balasubramanian and Shrivastava, 2020) by analyzing mapping failures and performing local adjustments to the schedule to obtain a shorter compilation time and better quality of the solution.
In our experiments, we compare our results to those obtained by both RAMP(Dave et al., 2018) and PathSeeker(Balasubramanian and Shrivastava, 2022). These two works had shown superior performance with respect to the earlier methodologies mentioned above, and therefore represent the current SoA of the modulo scheduling mapping problem. We quantitatively compare our work to them, and show that SAT-MapIt can better explore the scheduling space and get smaller II by using custom scheduling tables with a SAT formulation.
A second category of approaches addresses the mapping problem with Integer Linear Programming (ILP) or Boolean Satisfiability formulations. In (Chin and Anderson, 2018) the authors propose an ILP formulation approach and prove the feasibility of mapping in the given number of CGRA-cycles. Similarly, (Miyasaka et al., 2020) propose to use a SAT solver instead of an ILP solver to identify a valid solution. These works represent a first effort towards exact formulations, such as ours, but they are not capable of achieving a modulo scheduled solution. Our SAT formulation, initially introduced in (Tirelli et al., 2023), is to the best of our knowledge the first to propose an exact solution to the modulo scheduling problem, and yet scale to DFG sizes that were previously only tackled via heuristics.
Most SoA work lacks a direct link to the execution of mapped CILs on hardware. In fact, (Hamzeh et al., 2012; Chen and Mitra, 2014; Hamzeh et al., 2013; Oh et al., 2009; Park et al., 2008) all adopt abstract representations as architectural targets for their scheduling frameworks. This approach may lead to simplifying assumptions that do not reflect on hardware. As examples, (Chen and Mitra, 2014) and (Hamzeh et al., 2012) postulate that there are no constraints on instruction memory, while in (Hamzeh et al., 2012) and (Hamzeh et al., 2013) a large number of concurrent data memory transfers are allowed, without considering the capability of the system bus. An alternative approach is followed by the authors of (Dave et al., 2018) and (Wijerathne et al., 2022), who embed their CGRA compilation methodologies as part of system simulation frameworks, based on gem5 (Binkert et al., 2011) and SystemC, respectively. In this way, they ensure that the defined target hardware can be properly integrated in a system-on-chip. Nevertheless, system simulation waives the detailed description of the hardware implementation and hence does not offer a path to digital verification and synthesis, hampering the insights which can be gathered from an architectural perspective. Conversely, this works builds upon (Tirelli et al., 2023) to interface to an open-hardware CGRA implementation (Rodríguez Álvarez et al., 2023). In this way, we ensure the cycle-accurate correctness of SAT-MapIt mapping via post-synthesis simulations, reporting the performance and requirements of different CGRA sizes. Other works adopting this position are (Balasubramanian and Shrivastava, 2022) and (Chin and Anderson, 2018), where scheduled CILs are mapped on the CGRA-ME array (Chin et al., 2017). Of these, PathSeeker (Balasubramanian and Shrivastava, 2022) is most related to our work, since it also implements modulo scheduling. We comparatively evaluate our SAT-MapIt methodology against it in Section 5.
Figure 2. a): C code of our running example. The identified CIL is highlighted. b): LLVM IR of the identified CIL. c): Associated DFG.
Red edges are loop-carried dependencies, black edges are data dependencies.
3. Background
In this section, we provide the background needed to present our methodology, and we illustrate it wherever appropriate through a running example.
3.1. Compilation
To accelerate an application to a CGRA, a CIL is identified in the application, as in Figure 2.a; the identification can either be performed automatically through techniques such as, for example, (Zacharopoulos et al., 2018), or manually by the programmer through pragma annotations, as done in this work. Then, CIL must be compiled, in order to be translated into CGRA instructions. The first step in this process, depicted in Figure 2, is to generate a semantically equivalent version in Intermediate Representation (IR) (LLVM IR is our chosen one) and from there to DFG. DFGs are directed graphs in which nodes represent operations, edges represent dependency relations between operations, and loop-carried dependencies correspond to back-edges.
In a second phase, the generated DFG is mapped onto the CGRA, by assigning each of its nodes to a given PE in a given CGRA-cycle. To perform such mapping, a technique called modulo scheduling is employed and is described in the following.
3.2. Modulo Scheduling
Modulo scheduling is a compilation technique that enables efficient execution of a CIL, by executing multiple iterations of it in an interleaved manner. As exemplified in 2(a), a modulo-scheduled CIL is divided into three stages: prologue, kernel, and epilogue. Prologue and epilogue are one-time executed stages: the former is used to prepare the data to feed the pipeline, the latter to reorganize them at the end. The kernel is instead the steady state that executes multiple times and includes the operations to be parallelized through pipelining. The goal of modulo scheduling is to pipeline as effectively as possible the execution of CIL operations, and this corresponds to minimizing the II, i.e. the duration of the kernel stage – 3 CGRA-cycles in our example. Therefore, the mapping problem consists in finding a legal modulo schedule for a CIL, performing the placement and routing of operations in a constrained 3D space represented by the PEs dimensions and by time. An example of legal mapping for the DFG in our running example is shown in Figure 3.b.
Figure 3. a): Modulo scheduling of theDFG of the running example, highlighting the division between prologue, kernel, and epilogue.
b): Mapped CIL of the DFG in the running example on a $2\times 2$ CGRA
3.3. Problem Formulation
The mapping problem requires taking into account architectural constraints such as the number of available PEs, the available interconnections among PEs and memory, and the number of registers for each PE. Violation of such constraints cause invalid mappings leading to, e.g., the schedule of more operations than the number of available PEs during the same CGRA-cycle, or the impossibility to route a value between parent and children nodes. We formally define the problem as following:
Mapping Problem: Given a DFG $\mathcal{G(N,E)}$ representing a CIL, and given a CGRA architecture $\mathcal{A}$ in terms of size and topology, we generate a set of literals $\mathcal{L}$ in the form $v_{n,p,c,it}$ where $n$ is the node id, $p$ is the PE on which $n$ is mapped, $c$ is the CGRA-cycle at which $n$ is scheduled and $it$ is the iteration to which $n$ refers. We encode all constraints to which a mapping has to adhere to, in order to be valid, via a SAT formulation.
3.4. SAT problem
A Boolean satisfiability problem, namely (SAT), is the problem of determining if a Boolean formula is satisfiable or unsatisfiable. Usually, a problem is formulated in a Conjunctive Normal Form (CNF): a conjunction of clauses, each being a disjunction of literals.
For example, Equation 1 shows a CNF formula consisting of three literals $a$, $b$, and $c$, and three statements $a\vee b$, $b$, $c\wedge a$.
(1) $(a\vee b)\wedge b\wedge(c\wedge a)$
In order to determine the satisfiability of the CNF formula in Equation 1 a SAT solver searches for an assignment of the literals such that the formula is true. In this example, the answer is SAT, because setting a, b and c to true makes the formula evaluate to true.
In our work we must impose that only a fixed number of literals can be true at the same time. This constraint is called an at-most-K constraint and is well known in the literature(Bittner et al., 2019).
In the following, we detail the SAT formulation we devised to solve the mapping problem.
4. CGRA mapping Methodology
Our methodology addresses mapping by solving a SAT problem where data dependency, schedule and CGRA architecture are expressed as Boolean constraints in a CNF. Our toolchain takes as input the C code of the program, converts it into LLVM IR, extracts the designated CIL structures and, through a custom LLVM pass, obtains the information required to generate the DFG.
Figure 4. SAT-MapIt searches for mappings for a given II, iteratively increasing II in case the SAT solver returns UNSAT, or register allocation fails to color the model returned by the solver.
The information retrieved from the LLVM pass is then used to build a set of schedules, in turn translated into a set of constraints, which are finally fed to a SAT solver, as depicted in Figure 4. If the answer of the solver is SAT, the next step is to verify that there are enough registers available in the PEs to store the generated data for the whole liveness duration, and this is determined via Register Allocation (RA). If this further step also succeeds, then a valid mapping has been identified. Otherwise, the current II increases and the process is repeated iteratively.
Here, we detail the various steps of our methodology: schedule creation, SAT formulation, and finally register allocation.
4.1. Schedule creation
We first create As-Soon-As-Possible (ASAP) and As-Late-As-Possible (ALAP) schedules, for the input DFG. This corresponds to detecting how early and how late each node can be scheduled. We then generate the Mobility Schedule (MS), which expresses the mobility of each node from its $ASAP$ to its $ALAP$ position. Table 1 shows these schedules for our running example.
Table 1. ASAP, ALAP, and Mobility Schedule
Nodes
Time ASAP ALAP MS
0 1 2 3 4 3 1 2 3 4
1 5 7 10 4 5 1 2 4 5 7 10
2 6 11 1 6 7 1 2 4 5 7 10
3 8 2 8 10 1 2 4 5 7 10
4 9 9 11 9 11
At this point, we create the Kernel Mobility Schedule (KMS): an ad hoc structure used to formulate the mapping problem with SAT-MapIt. The KMS can be seen as a superset of all possible mappings, and is the product of iteratively folding MS by an amount equal to II: every time MS is folded by II into KMS, each node receives a label that refers to the iteration number it belongs to. The more folding we have, the more iterations our CIL will execute in the kernel. This process is depicted in Table 2, again for our running example. Given an II of 3, the MS is folded twice ($\lceil 5/3\rceil=2$) and, hence, the KMS contains two iterations. The first is depicted in blue, and the second in green.
Table 2. CIL Mobility Schedule creation.
In blue iteration 0, in green iteration 1
MS
Time Nodes
0 1 2 3 4
1 1 2 4 5 7 10
2 1 2 6 7 10 11
3 2 8 10 11
4 9 11
KMS
Time Nodes
0 1 2 6 7 10 11
1 2 8 10 11 1 2 3 4
2 9 11 1 2 4 5 7 10
Together, DFG and MKS are used to generate all the statements of the CNF formulation of the mapping problem, which is the subject of the next subsection. The iterative search for a valid mapping can be shortened by considering that available resources and loop carried dependencies impose a lower bound on the achievable II. Such minimum II (mII) as defined in (Rau, 1996), is hence expressed as
(2) ${\leavevmode mII}=max(ResII,RecII)$
[…]
4.2. SAT formulation
We create a CNF formula using literals in the following form: $x_{n,p,c,it}$ , where $n$ denotes the node identifier in the DFG, $p$ denotes a PE on the CGRA, $c$ represents at which CGRA-cycle a node is scheduled, and $it$ to which iteration the node refers to. For example, the literal $x_{3,2,1,0}$ represents whether node 3 is mapped on PE2, at time 1 and iteration 0.
Our problem formulation can be described at a high level by partitioning all statements into three main sets of clauses that assure the following:
  • C1: Every node is associated with a set of literals, and for each one of those sets, one and only literal must be set to True
  • C2: At most one node should be assigned to a PE at a given CGRA-cycle, since two or more nodes cannot be scheduled simultaneously on the same PE.
  • C3: Each node’s predecessor and/or successor must be assigned to a neighbor or on the same PE.
To provide a formal description of the above constraints, we introduce some additional definitions. Let $\mathcal{L}$ be the set of all literals, then $\mathcal{L}(n)$ be the set of all literals associated with node $n$. For example, for node $n_{3}$ of the running example we would have:
(3) $\mathcal{L}(n_{3})=\{x_{3,0,1,1},x_{3,1,1,1},x_{3,2,1,1},x_{3,3,1,1}\}$
because, in the KMS, in Table 2 node 3 appears only at time 1 and at iteration 1, and can be mapped in any PE 0,1,2,3.
To make the notation more compact and easy to read, we also associate each literal in the form $x_{n,p,c,it}$ to a literal written as $v_{i}$.
So, Equation 3 can also be written as:
$\mathcal{L}(n_{3})=\{v_{0},v_{1},v_{2},v_{3}\}$
where $v_{0}=x_{3,0,1,1}$, $v_{1}=x_{3,1,1,1}$, $v_{2}=x_{3,2,1,1}$ and $v_{3}=x_{3,3,1,1}$.
Now, we can start the description of the three sets of constraints. The first set of constraints, C1, ensures that all the nodes are mapped on the CGRA, and can be encoded formally with:
(4) $\phi(n)$ $=\bigvee_{v_{i}\in\mathcal{L}(n)}v_{i}$
$\alpha(n)$ $=\bigwedge_{(v_{i},v_{j})\in\mathcal{M}(n)}\neg(v_{i}\wedge v_{j})$
$\beta(n)$ $=\phi(n)\wedge\alpha(n)$
where $n$ is one of the nodes in the DFG and $\mathcal{M}(n)$ is defined as follows:
$\mathcal{M}(n)=\{(v_{i},v_{j}):v_{i}\prec v_{j},(v_{i},v_{j})\in\mathcal{L}(n)\times\mathcal{L}(n)\}$
where $v_{i}=x_{n,p_{1},c_{1},it_{1}}$ and $v_{j}=x_{n,p_{2},c_{2},it_{2}}$ with $p_{1}\neq p_{2}$, $c_{1}\neq c_{2}$ and $it_{1}\neq it_{2}$. Furthermore the symbol $\prec$ represents the lexicographically smaller-than relation between two literals. For example $x_{3,0,1,0}$ is lexicographically smaller than $x_{3,1,0,0}$, while $x_{3,1,0,1}$ is not.
Equation 4 will be used on each node of the DFG and each $\beta$ generated will be added to the SAT solver. In our example with $n_{3}$ Equation 4 becomes:
$\phi(n_{3})$ $=v_{0}\vee v_{1}\vee v_{2}\vee v_{3}$
$\alpha(n_{3})$ =$=\neg(v_{0}\wedge v_{1})\wedge\neg(v_{0}\wedge v_{2})\wedge\neg(v_{0}\wedge v_{3})\wedge$
$\neg(v_{1}\wedge v_{2})\wedge\neg(v_{1}\wedge v_{3})\wedge\neg(v_{2}\wedge v_{3})$
$\beta(n_{3})$ $=\phi(n_{3})\wedge\alpha(n_{3})$
The second set of constraints, C2, avoids solutions that map on the same PE more than a single nodes at the same time, and is encoded through:
(5) $M(n,m)$ $=\bigwedge_{(v_{i},w_{j})\in\mathcal{V}(n,m)}\neg(v_{i}\wedge w_{j})$
$\xi$ $=\bigwedge_{n}^{N-1}\bigwedge_{m=n+1}^{N}M(n,m)$
with $\mathcal{V}(n,m)$ defined as:
$\mathcal{V}(n,m)=\{(v_{i},w_{j}):v_{i}\prec w_{j},v_{i}\in\mathcal{L}(n),w_{j}\in\mathcal{L}(m)\}$
The last set of constraints, C3, handles the dependencies in the DFG, and requires a lengthier explanation. For each dependency, we consider only literals that are at most one iteration apart in the KMS, on a neighbor PE and that respect one of the relations:
(6) $\left\{\begin{array}[]{@{}ll@{}}c_{d}\leq c_{s}\text{ if }it_{s}\neq it_{d}\\ c_{d}>c_{s}\text{ if }it_{s}=it_{d}\end{array}\right.$
where $c_{d}$ is the CGRA-cycle at which the destination node is scheduled, and $c_{s}$ is the CGRA-cycle at which the source node is scheduled. For example, note that for $n_{s}=n_{10}$ and $n_{d}=n_{11}$, at $c_{s}=1$ and $c_{d}=2$ at $it_{s}=0$ and $it_{d}=0$ the second relation of Equation 6 is satisfied. This constraint ensures that a node consumes the value produced by the predecessor in the proper order, avoiding overlapping of the same dependencies among kernel and prologue/epilogue stages.
Before proceeding in the description of C3, we define two functions that will be used to better formalize the constraints. The neighborhood function expresses whether the PEs of two literals are neighbors, and is defined as follows:
(7) $f_{n}(v_{1},v_{2})=\left\{\begin{array}[]{@{}ll@{}}2&\text{if neighbors and}\ p_{1}\neq p_{2}\\ 1&\text{if neighbors and}\ p_{1}=p_{2}\\ 0&\text{otherwise}\end{array}\right.$
The CGRA-cycle function returns the CGRA-cycle associated to a literal, e.g. $f_{c}(v_{i})=f_{c}(x_{n,p,c,it})=c$.
C3 enforces that each data dependency is mapped on neighbors PE on the CGRA and that data produced by a node on a CGRA-cycle is consumed before being overwritten by another node on a subsequent CGRA-cycle. For each edge in the DFG, we generate a set of constraints $\xi$ that will be added to the SAT solver, and then we take into account all the possible combinations of the same dependency in the KMS:
[…]
Now we proceed to give all the information needed to formally define C3:
(9) $\mathcal{D}(n_{s},n_{d})=\left\{\begin{array}[]{@{}ll@{}}\gamma&\text{if}\ distance=1\\ \zeta&\text{if}\ distance\neq 1\\ \end{array}\right.$
where the $\mathcal{D}(n_{s},n_{d})$ refers to the constraint associated to the edge $(n_{s},n_{d})$ that depending on the distance value can be either $\gamma$ or $\zeta$. The distance between two nodes in the KMS is computed with the following equation:
[…]
Figure 5. Top: two examples of dependencies – $S_{1}$, $D_{1}$ and $S_{2}$,$D_{2}$ – with a distance equal to 1.
Bottom: an example of dependence with a distance greater than 1.
For example, let us consider $n_{s}=n_{2}$ of iteration 0 at CGRA-cycle 0 ($c_{s}=0$ and $n_{d}=n_{9}$ of iteration 0 at CGRA-cycle 2 ($c_{d}=2$). In this case, we have $(2-0+3)\mod 3=2$.
Equation 10 returns the number of CGRA-cycles that the source node needs to wait for, until its value is consumed. We use this information to distinguish between two main cases depicted in Figure 5. The figure on the top shows the case when source and destination node, on different PEs, are only one CGRA-cycle apart; on the bottom, source and destination are more than one CGRA-cycle apart.
We now proceed to define $\gamma$ and $\zeta$.
(11) $\gamma=\bigvee_{(v_{i},w_{j})\in\mathcal{A}(n_{s},n_{d})}(v_{i}\wedge w_{j})$
where $\mathcal{A}(n_{s},n_{d})$ is a set of literals defined as:
(12) $\mathcal{A}(n_{s},n_{d})=\{(v_{i},w_{j}):f_{n}(v_{i},w_{j})>0,v_{i}\in\mathcal{L}(n_{s}),w_{j}\in\mathcal{L}(n_{d})\}$
Now we define $\zeta$ as the or between two sets of literals.
(13) $\zeta=\zeta_{1}\vee\zeta_{2}$
where $\zeta_{1}$ is defined as:
(14) $\zeta_{1}=\bigvee_{(v_{i},w_{j})\in\mathcal{B}(n_{s},n_{d})}(v_{i}\wedge w_{j})$
and where $\mathcal{B}(n_{s},n_{d})$ is a set of literals defined as:
(15) $\mathcal{B}(n_{s},n_{d})=\{(v_{i},w_{j}):f_{n}(v_{i},w_{j})=1,v_{i}\in\mathcal{L}(n_{s}),w_{j}\in\mathcal{L}(n_{d})\}$
Lastly we define $\zeta_{2}$ as:
(16) $\zeta_{2}=\bigvee_{(v_{i},w_{j})\in\mathcal{A}(n_{s},n_{d})}\left(v_{i}\wedge w_{j}\wedge\neg\left(\bigvee_{z_{k}\in\mathcal{V}(ns,n_{d})}z_{k}\right)\right)$
where $\mathcal{A}(n_{s},n_{d})$ is constructed as in Equation 12 and $\mathcal{V}(n_{s},n_{d})$ is the set of literals with the same PE of the source node on different CGRA-cycle defined as:
(17) $\mathcal{V}(n_{s},n_{d})=\{z_{k}:f_{c}(v_{i})<f_{c}(z_{k})<f_{c}(w_{j}),z_{k}\in\mathcal{L}\}$
The CNF $\zeta$ is composed of two terms because $\zeta_{1}$ handles the case in which the output of the source node is delivered to the destination node through the registers on the PE ($n_{s}$ and $n_{d}$ necessary on the same PE), while $\zeta_{2}$ handles the case in which the output is written in the output register of the PE, and hence it must not be overwritten in subsequent CGRA-cycles.
Figure 6. Top: the PE used by the source node must be left empty until the destination node consumes its value.
Bottom: the PE used by the source node can be also used in subsequent CGRA-cycles, as the value written by the source node is, in this case, saved in a register (this will require Register Allocation to be validated).
Figure 6 shows those two cases. The figure on top shows that the PE 0 at CGRA-cycle 1 cannot be filled with another instruction, since it will overwrite the data produced by $n_{s}$. The figure on the bottom shows instead that the PE 0 at CGRA-cycle 1 can be overwritten since the destination node is on the same PE as the source node and an internal register can be used. The SAT solver will then decide which of the two possibilities is used to generate a mapping. Giving both these options to the solver provided the key to find better mapping in several cases, as mentioned in the Experimental Section. Lastly, back dependencies need to be handled in a slightly different way. In particular, Equation 6 becomes:
(18) $\left\{\begin{array}[]{@{}ll@{}}c_{d}\leq c_{s}\text{ if }it_{s}=it_{d}\\ c_{d}>c_{s}\text{ if }it_{s}>it_{d}\end{array}\right.$
This concludes the description of the third set of constraints. For the running example, the satisfying solution given by the SAT solver sets the following literals to true: $v_{11,1,0,0}$, $v_{6,2,0,0}$, $v_{7,3,0,0}$, $v_{2,0,1,0}$, $v_{1,1,1,1}$, $v_{8,2,1,0}$, $v_{3,3,1,1}$, $v_{9,0,2,0}$, $v_{10,1,2,1}$, $v_{4,2,2,1}$, $v_{5,3,2,1}$. This, in turn, corresponds to the mapping shown in Figure 3.b.
Register Allocation is the last phase needed in order to validate the mapping returned by the solver.
4.3. Register Allocation
In our CFN formulation, to have a more straightforward set of constraints, we chose not to include information about which register should store the data produced by the node on every PE. We hence treat the RA task as a separate problem.
After the solver returns a mapping, the RA phase needs to ensure that the CGRA has an adequate number of registers to execute the CIL correctly. SAT-MapIt  as a separate phase subsequent to SAT solving, exploits the SSA format of the input code and looks for an optimal solution(Hack et al., 2006) of the RA problem.
For each PE in the CGRA, we construct an interference graph that we then attempt to color using $n$ colors, where $n$ is the number of registers available on each PE. If the coloring succeeds, we know that no further action is needed, and the toolchain provides the assembly code for the CGRA. However, if the coloring process fails, we increment the II and initiate a new search. While it’s possible to explore additional graph splitting techniques to make the graph colorable, we do not consider them herein, as such optimizations are orthogonal to the mapping problem we address in this work.
5. Experimental Results: SAT-MapIt solutions
In this section we evaluate the effectiveness of SAT-MapIt on a set of CILs from MiBench and Rodinia benchmark suites, shown in Table 3. We compare the obtained II, and the time to find it, with respect to two techniques of the SoA: RAMP(Dave et al., 2018) and PathSeeker (Balasubramanian and Shrivastava, 2022).
5.1. Experimental Setup
We use the original code publicly released by the authors of the two SoA tools and the same DFG, while in performance evaluation we extract the DFG directly from the C code and apply an instruction selection pass to convert the LLVM-IR instructions into instructions compatible with the ISA of the CGRA used. In the target CGRA architecture that we consider in our experiments, each PE is connected to the four nearest neighbors, as in Figure 9, and each PE contains four local registers. We vary the size of the mesh from $2\times 2$ up to $5\times 5$. The Z3 solver (Moura and Bjørner, 2008) is used to solve our SAT formulation. All experiments are performed on a machine with 2.6 GHz 6-Core Intel Core i7. For PathSeeker, each experiment was repeated 10 times.
Table 3. List of benchmarks with graph information
Suite CIL #nodes #edges
MiBench sha 30 33
sha2 26 28
gsm 20 24
patricia 42 46
bitcount 26 29
basicmath 19 20
stringsearch 16 16
Rodinia backpropagation 35 39
nw 16 16
srand 22 22
hotspot 67 76
5.2. SAT-MapIt achieves better II
The performance of a mapping is first and foremost measured by the II achieved, because this, in turn, is a measure of the level of parallelism obtained; our first experiments hence compare the II of SAT-MapIt with those of the SoA, for each benchmark explored. This is depicted in Figure 7, which shows the performance obtained by all techniques for different CGRA sizes. For SoA, we report the best result returned by the two algorithms.
Figure 7. Experimental results of the chosen benchmarks for different architecture sizes. We compare the II found by SAT-MapIt with respect to the best results obtained by RAMP and PathSeeker – lower is better. A red cross means that the process did not terminate before a timeout of 4000 seconds. A black cross means that the process was terminated when it reached a current II of 50 but still found no feasible solution. Red dashes indicate the minimum II (mII) for that CIL. For the $2\times 2$ CGRA, hotspot had a mII of 17, which is not displayed.
It can be seen from Figure 7 that SAT-MapIt sets a new benchmark in the field as it almost always either outperforms the SoA or achieves the II lower bound (mII), for the evaluated CILs. In particular, SAT-MapIt found a solution in 6 out of the 7 cases where the SoA could not; it reached a lower II than the other tools for 14 CILs and reached the mII (computed from Equation 2) for 34 CILs while the SoA could only achieve so for 19 of them. This indicates that SAT-MapIt  explores the solution space more effectively than SoA techniques, and this effect is especially visible when the mapping problem becomes more challenging, as detailed in Section 5.3. The three exceptions to these patterns are hotspot in the $2\times 2$ CGRA, where no tool could find a solution for an mII of 19 before the timeout of 4000 seconds; backprop in the $4\times 4$ CGRA, where SAT-MapIt matches the SoA but without reaching the mII; and sha in the $5\times 5$ CGRA. We further comment on this last result in Section 5.5.
5.3. SAT-MapIt uses tight resources better
By focusing on the $2\times 2$ size CGRA, it can be seen not only that SAT-MapIt  outperforms SoA in 8 of 11 benchmarks, but also that it can find a valid solution three times (sha, patricia and backpro) where SoA could not. This showcases the effectiveness of our methodology, particularly when the mapping problem becomes more challenging, always managing to use the minimal number of resources possible. The other tools evaluated either return no solution, or seem to be biased towards adding routing nodes even when they are not necessary in order to find a solution. Our tool can then fully exploit and handle the resources available, even when these are tightly constrained e.g. in low-power edge architectures (Duch et al., 2019).
Table 4. Mapping time (seconds) for different CGRA sizes
$2\times 2$ $3\times 3$ $4\times 4$ $5\times 5$
CIL SoA Ours $\Delta$ SoA Ours $\Delta$ SoA Ours $\Delta$ SoA Ours $\Delta$
sha 41.01 3.22 -37.79 0.23 2.86 2.63 32.04 7.23 -24.81 6.25 28.93 22.68
gsm 2.81 1.25 -1.56 4.14 4.35 0.21 32.04 10.46 -21.58 0.29 21.4 21.11
patricia 1351 5.39 -1345 48.98 16.31 -32.67 421.05 39.28 -381.77 4000 75.16 -3924
bitcount 2.63 1.68 -0.95 5.86 7.84 1.98 0.44 21.55 21.11 0.01 47.62 47.61
backprop 1262 3.39 -1259 44.62 12.27 -32.35 57.17 25.1 -32.07 981.73 52.43 -929.3
nw 0.01 0.56 0.55 0.03 1.56 1.53 0.08 3.63 3.55 0.02 7.75 7.73
srand 0.32 1.15 0.83 0.09 3.9 3.81 0.25 8.79 8.54 0.02 21.64 21.62
hotspot 4000 4000 0 13.19 28.43 15.24 3556.62 3734 178.15 4000 108.02 -3891
sha2 4000 2.21 -3997 61.7 3.13 -58.57 696.6 7.19 -689.41 675.12 16.88 -658.24
basicmath 0.01 0.62 0.61 0.07 1.9 1.83 0.22 4.11 3.89 0.5 8.7 8.2
stringsearch 0.19 1.02 0.83 3.27 3.55 0.28 0.02 7.62 7.6 0.02 15.3 15.28
5.4. SAT-MapIt is faster when run-times are high
Given that the II found are better than SoA, it may be interesting to also analyze the time needed to find such solutions; this is reported in Table 4. The following statements can be made when looking at these numbers. 1) Our tools running time is longer than SoA in 26 out of 44 experiments. 2) However, in these 26 cases the average time difference is only 15.28 seconds, with a standard deviation of $4.973$) On the other hand, in the 18 cases in which our tool is faster, the average time difference is 962.24 seconds, with a standard deviation of $1438.78$. This shows that SAT-MapIt is significantly faster when it matters, i.e. when computation times are high.
5.5. Limitations of SAT-MapIt
Currently, our tool does not apply any routing strategies. This limitation manifests in the sha CIL of a $5\times 5$ CGRA, where we achieve an II of 3, while (Dave et al., 2018) and (Balasubramanian and Shrivastava, 2022) can find an II of 2 by adding a routing node. This is the only case, out of the 44 experiments shown, where the effect of this limitation can be noticed. Implementing routing capabilities in SAT-MapIt will be the subject of future work.
5.6. An example of mapping
Figure 8. Mapping results of sha1 CIL for both tools. RAMP on the right, SAT-MapIt  on the left
Lastly, we show in more detail one of the mappings where we achieved a better II than SoA: benchmark sha1 and CGRA size $3\times 3$. Figure 8 shows the two mappings obtained for this benchmark: to the left, by SAT-MapIt, and to the right by RAMP. The DFG contains 30 nodes and 33 edges – all nodes and only a subset of the edges are represented in the picture. In the description of our methodology, and in particular in how to handle dependencies, we showed in Figure 6 how SAT-MapIt can handle two different cases of mappings. This capability will be showcased here, considering in particular a dependence from node 23 to node 0, and then another one from node 12 to node 13. We can see that the PE3 is left empty until the data is consumed, which is one of the options available to the SAT solver and encoded through Equation 18. Instead, consider now the dependence between node 12 and node 13. In that case, we can see that the constraint associated with it is the one in Equation 14, which allows node 8 to be mapped on the same PE of node 12, exploiting the internal registers of PE 4. With these capabilities, SAT-MapIt finds a lower II, specifically ${\leavevmode II}=4$. While RAMP returns a mapping with ${\leavevmode II}=5$. We can also notice that node 46 (in red) is an unnecessary routing node that RAMP added to resolve a dependency and map the CIL.
6. Application-to-hardware framework
Figure 9. a): Top-level view of the openEdgeCGRA architecture with a $4\times 4$ PE array, b): PE-level architectural view
SAT-MapIt operates on an abstract view of applications and architectures. In particular, it relies on an instructions-level representation of CILs, according to the single-assignment form provided by the LLVM IR. Such hardware-agnostic stance allows to flexibly target any CGRA architecture supporting modulo scheduling ISA (Duch et al., 2019; Singh et al., 2000). Nonetheless, hardware-specific back-ends must be employed to translate the compiler output to a binary representation compatible with a target architecture. After back-end translation, the resulting bitstream contains a sequence of control words, dictating the desired behavior of the PEs at CIL execution time, for each CGRA-cycle. These instructions encode the operation performed by ALUs, the input operands, and the output destination. Note that, in modulo scheduled CGRAs, several instructions are configured in each PE, with one being executed in all PEs at every CGRA-cycle, to implement modulo scheduling as discussed in Section 3. Individual bits of instructions directly or indirectly drive the multiplexers that are responsible for the control logic.
A direct connection between the control words and multiplexers is featured in openEdgeCGRA (et al., 2020), the CGRA considered in this work, waving the need for a decoder for faster CIL execution. In particular, we target the implementation available in the public repository in  (Rodríguez Álvarez et al., 2023). It is fully synthesized and can be simulated at the post-synthesis level, allowing us to inspect the performance and energy requirement of mapped CILs, which are presented in Section 5.
The openEdgeCGRA architecture, shown in Figure 9, comprises a two-dimensional array of PEs, interconnected with their nearest neighbors in a torus, wrapping around columns and rows (Figure 9.a). PEs are organized into columns, each having its own Program Counter (PC). At the array periphery, a Direct Memory Access (DMA) engine interfaces PEs with the system bus, providing one input/output port per CGRA column. A context memory acts as a cache for configurations, storing the bitstreams encoding the required CILs. The descriptor of the CIL (index, required number of columns and location of the configuration words in the context memory) are stored in the Kernel Memory. Each PE is composed of data registers, a program memory and an ALU (Figure 9.b). The program memory governs the local behavior of the PE according to the value of the word indexed by the PC. Each program memory word is composed of 32 bits, and encodes the source of operands (immediate, register file, neighboring cells or main memory via the system bus), where the output is stored, and the operation being performed.
ALUs can perform arithmetic operations (signed addition, subtraction and multiplication, as well as fixed-point multiplication), arithmetic and logic shifts, and bit-wise operations. The ALUs also supports selecting operands based on the state of the zero and sign flags (BXFA and BSFA, respectively), thus allowing the implementation of branches via if-conversion. Data transfers to/from memory are realized by load and store opcodes. Finally, PEs can conditionally or unconditionally overwrite the value of program counters by issuing branch and jump instructions. Such instructions are used to govern mapped modulo-scheduled CILs, so that the prologue is executed once, the kernel is iterated as required by the loop count, before ending with the epilogue. Local PE data storage is provided by a 4-word register file, the PE output register, and a flag register storing the 1-bit zero and sign flags resulting from the previous instruction.
The CGRA ISA closely mimics the integer set of opcodes supported by LLVM, as reported Table 5. This characteristic results in a simple back-end implementation. On the other side of the coin, it restricts its support to integer CILs. Hence, we only considered integer benchmarks for the exploration results discussed in Section 7.
Table 5. Instructions set architecture of the open hardware CGRA ISA, as reported in (Rodríguez Álvarez et al., 2023).
Type of instruction Opcode
 Arithmetic operations SADD, SSUB, SMUL, FXPMUL
 Shifts SLT, SRT, SRA
 Bit-wise operations LAND, LOR, LXOR, LNAND, LNOR, LXNOR
 Selects BSFA, BZFA
 Loads and stores LWD, LWI, SWD, SWI
 Conditional and unconditional branches BEQ, BNE, BLT, BGE, JUMP
 No operation NOP
 Finish EXIT
At run-time, when a CIL execution is requested, its descriptor is fetched from the Kernel Memory, and the synchronizer module checks if enough PE columns are idle. If this is not the case, it waits for enough resources to become available, otherwise it copies the configuration words from the context memory to the program memory of PEs. CILs are then executed according to the selected configurations, governed by the value of program counters. Execution ends when a PE issues the exit instruction.
7. Experimental results: run-time performance of compute-intensive loops mapped with SAT-MapIt
In this Section, we investigate the characteristics of mapped CILs from a hardware perspective, showcasing the effect of provisioning different amounts of resources on performance and energy efficiency. Bridging compiler-level and hardware-level explorations, we then inspect the degree of correlation between respective metrics in the two spaces: Iteration Interval and Utilization in the former case; Energy, Latency, in the latter one.
7.1. Experimental Setup
We consider square openEdgeCGRA instances with sizes $2\times 2$, $3\times 3$ and $4\times 4$, referred to as $D2$, $D3$ and $D4$, respectively, in the following. Benchmark CILs from the MiBench suite were executed on the architectures. To compare with the SoA in Section 5.2 we used the same DFGs generated by the RAMP toolchain; however those DFGs are not inherently compatible with the target CGRA’s ISA, preventing the direct execution of experiments using identical graphs. To address this discrepancy, we recalculated the DFGs for each CIL to align with the instruction set of openEdgeCGRA, for this part of the experiments. Consequently, this resulted in different II values compared to Figure 7.
Because execution time can vary across runs based on input data, all inputs were provided from a 32-bit pseudo-random number generator adapted from (L’Ecuyer, 1999). Furthermore, where the array lengths could vary in each run, these were fixed as reported in Table 6. Across sizes, CILs received the same series of inputs to allow a fair comparison.
[…]
Table 6. Compiler level (II, U) and run-time level (Energy, Latency) performance of benchmark CILs.
II U (%) Energy (nJ) Latency (clock cycles)
CIL $D2$ $D3$ $D4$ $D2$ $D3$ $D4$ $D2$ $D3$ $D4$ $D2$ $D3$ $D4$
reversebits 3 3 3 75 33 19 1.5 2.1 2.8 155 158 159
bitcount 4 4 3 38 17 15 1.0 1.3 1.6 121 122 110
sqrt * 5 5 5 40 18 10 1.5 1.9 2.7 174 174 176
stringsearch 4 2 2 100 89 50 13.5 5.1 6.5 1,150 349 351
gsm 5 3 4 70 52 22 4.2 4.7 7.0 389 326 388
sha 7 4 3 89 69 54 13.5 13.4 14.4 1,150 862 653
sha2 8 8 36 20 4.7 6.0 289 289
    Input values are limited to unsigned 31 bits.
  • Input values are limited to [1, 255]. Pattern length patlen is set to 50 words. Variable skip2 is fixed to an arbitrary constant value.
  • Input values are limited to signed 16 bits. Array length is set to 40 words.
7.2. Experimental Results
Obtained run-time latency and energy values across benchmarks and architectures are summarized in Table 6. In addition, the table also reports compiler-level metrics: II and U. Experimental results are discussed in the following, providing insights on the inter-relationship among architectural parameters (CGRA size), compiler-level metrics and run-time ones.
Relationship between energy and size.
It can be observed that smaller CGRAs are more energy efficient in those cases in which increasing the CGRA size does not result in a decrease in II, as seen in Table 6 for CILs reversebits and sqrt for all sizes, where the $D2$ CGRA has the smallest energy envelope. This amplifies the importance of a compiler that is able to perform scheduling not just when real estate in the CGRA s is abundant, but even when the resource budget is tight, such as the compiler presented here: the most important energy savings are achieved in this case.
Relationship between II and size
In almost all CILs, II has a monotonic relationship with CGRA size. The gsm CIL is an exception where $D3$ has the best II. In openEdgeCGRA, this counter-intuitive scenario is found when the mapping exploits the wrap-around between the PEs of the edge. When the CGRA size increases, this connection is lost and a less efficient mapping is achieved. If nodes in a section of the DFG are fully connected (all nodes are related to each other), they can be efficiently mapped in a set of PEs that are also fully connected. Figure 10.a illustrates how the $D3$ CGRA offers the highest number of fully connected PEs. Therefore, if the sample DFG from Figure 10.b is mapped into the three CGRA dimensions, the best II that can be obtained with SAT-MapIt is in $D3$. Such mapping is shown in Figure 10.c along with its counterparts for $D2$ and $D4$.
Figure 10. a: In a 2-PEs and 3-PEs row (or column), processing elements are fully connected in a torus mesh. Adding an extra PE reduces the number of fully connected PEs. b: An example DFG where full connectivity leads to a lower II. c: The mapping of the DFG from b) across different CGRA dimensions. Each row represents a CGRA-cycle on one row of PEs. An II of 2 cycles is only achievable if 3 fully connected PEs are available.
Relationship between II and latency.
As expected, II and latency have a monotonic relationship. However, changes in II and latency are not proportional as the latter metric also captures architecture-specific characteristics. In particular, II does not consider the different delay of instructions in the ISA. In openEdgeCGRA, load instructions from Table 5 take 2 clock cycles, while other instructions require a single one. Therefore, because the CGRA’s execution of an instruction is blocked by its slowest PE, packing all slow instructions in the same CGRA-cycle might yield a faster execution than spreading them throughout the CIL. In addition, timing is affected when access to shared resources is serialized by the hardware. In our target architecture, concurrent store instructions to the same memory bank are pipelined. Furthermore, read instructions in the same column (even aiming at different memory banks) are also serialized. The arbitration among memory accesses hence increases run time, as the execution of CGRA instructions must be paused until all memory accesses scheduled concurrently terminate. These effects can be observed in Table 6 for stringsearch going from $D3$ to $D2$: the benchmarks can be mapped in the smaller architecture by doubling the II, but such penalty is exacerbated by resources contentions, leading to a tripling of run-time latency.
Figure 11. a), d): Performance-efficiency solution space of mapped CILs, in terms of compiler-level metrics (a) and run-time metrics (d). b), c): relation between performance and efficiency metrics in the two spaces. Data is reported for each CIL across CGRA sizes and is normalized per CIL. Overlaps are slightly offset for visualization.
Relationship between compiler-level and run-time metrics
While the II provides clues about execution latency, similarly U (the ratio of non-idle PEs across all mapped instructions) gives hints about the power efficiency of a mapping. In this light, Figure 11.a illustrates the performance/efficiency trade-offs exposed by different CGRA sizes for the considered benchmarks using these compiler metrics [^1] For clarity, the graph reports under-utilization (Under-U), that is, the ratio of idle PEs over all mapped instructions, which is positively correlated to energy requirements.. Figure 11.d instead reports the same trade-off space between performance and efficiency expressed with run-time metrics (energy and latency) from post-synthesis simulations. In all cases, values in Figure 11 are normalized per CIL to the maximum among the three different sizes. Figure 11.b and 11.c plots the relation between corresponding metrics in the compiler-level and run-time space: II vs. Latency, Energy vs. Under-U, respectively. These intermediate plots highlight two trends. In Figure 11.b, a decrease in II caused by the size increase reflects in an almost linear drop in latency. In Figure 11.c the increase in U caused by the size decrease is reflected in a drop in energy consumption limited by a fixed cost (which is seen as a curved down-slope).
However, there are outliers to these patterns. In particular, the gsm benchmark has a minimum latency for the $D3$ architecture, as mentioned above. Moreover, the stringsearch overall energy consumption almost triples when going from $D3$ to $D2$ due to the drastic increase in latency with II. This effect cannot be perceived in Figure 11.a, but becomes evident in subplots b, c, and d, where implementation details are considered. These examples show that considering mapping results from a purely II (Karunaratne et al., 2017; Mei et al., 2007; Hamzeh et al., 2013; Dave et al., 2018; Friedman et al., 2009; Balasubramanian and Shrivastava, 2022) or U (Wijerathne et al., 2021; Mei et al., 2007) perspective does not completely capture the performance of mapped CILs. Nonetheless, they can still be effectively employed to prune the design space of possible implementations, in advance of time-consuming detailed hardware simulations. In particular, the Pareto sets of best performing implementations (i.e., the ones which are not dominated in terms of performance/efficiency) is highly similar in Figure 11.a and Figure 11.d. Considering all benchmarks, 83% of the Pareto points in the (II, U) space are also Pareto points in the (energy,latency) space. Dually, all Pareto points in Figure 11.d are part of the Pareto set in Figure 11.d. Hence, only selecting Pareto solutions in the architecture-agnostic (II, U) space for detailed (but lengthy and architecture-specific) hardware simulation does not result in sub-optimal outcomes, while pruning the exploration space by 40%.
8. Conclusion
In this paper, we have presented a tool, called SAT-MapIt, for modulo scheduling CILs onto CGRAs. We found that previous techniques, mainly based on classic graph algorithms such as Max-Clique enumeration, do not always explore the scheduling space effectively. Therefore, we have proposed a new SAT formulation of the modulo scheduling problem on CGRA that fully explores the scheduling space and finds the lowest II possible for a given DFG. To define the mapping problem through a SAT formulation, we also introduce a new ad-hoc schedule called Kernel Mobility Schedule, which is used with the DFG of the CIL to be mapped, and with the architectural information of the CGRA, to generate all the constraints that the SAT solver needs to obey. Overall, SAT-MapItfinds better solutions with respect to the SoA alternatives(Dave et al., 2018; Balasubramanian and Shrivastava, 2022), achieves better results in around 50% of the cases, and even identifies valid mappings where other tools could not find a valid solution. Compiled CILs where mapped on openEdgeCGRA in  (et al., 2020). We show how the evaluation of architecture-agnostic compiler-level metrics can effectively guide the mapping process, even if it cannot provide a complete view of run-time efficiency and performance. Moreover, we demonstrate how it can be employed to successfully prune the design space in a fast way, in advance of laborious hardware explorations.